5. The equilibrium concentrations of the reacting species in the reaction below are CO)-0.50M: [Hs)-0.45M: [CO2]-0.86M and (H20] 0.40 M a calculate Ka for the reaction at 5000 b, if we decrease the CO, concentration to 0.5M; what will the concentrations of all the gases be when equilibrium is reestablished? c. what is the Kp of this system? CO2 (g) + H2 (g) ←→CO (g) + H,O(g) 6. You are conducting a titration of 22 ml ofo.1 M HCI with 0 1 M NaOH How many mi will it take to reach the equivalence point?

media%2Fa4f%2Fa4f8c47b-d23c-4575-a56a-01
Answer #1

7) pH = pka + log(ch3coona/ch3cooh)

no of mol of ch3cooH = 40*0.1 = 4 mmol

no of mol of ch3cooNa = 10*0.2 = 2 mmol

pka of ch3cooh = 4.74

pH = 4.74+log(2/4)

   = 4.44

8) [H+] = 1.9*10^-3 M

   pH = -log(1.9*10^-3)

      = 2.72

   answer: a.acidic

9) answer: d.all of these.

10. answer: b.bronsted acids.